OpenStudy (acacia):

Solve for x. 6(x – 1) = 9(x + 2) x = –8 x = –3 x = 3 x = 8

KjSaif (kjsaif):

Let's solve this equation step-by-step. :) @acacia

KjSaif (kjsaif):

Simplify both sides of the equation.

OpenStudy (acacia):

Ok.. done

KjSaif (kjsaif):

(6)(x)+(6)(−1)=(9)(x)+(9)(2)(Distribute) 6x+ −6 = 9x+18 6x−6= 9x+18

KjSaif (kjsaif):

6x−6= 9x+18 is this what you got? if you did then good job! :)

KjSaif (kjsaif):

Subtract 9x from both sides.

OpenStudy (akgorgeous):

-8

KjSaif (kjsaif):

6x−6−9x=9x+18−9x −3x−6=18

KjSaif (kjsaif):

Add 6 to both sides.

KjSaif (kjsaif):

−3x−6+6=18+6 −3x=24

OpenStudy (acacia):

Ok

KjSaif (kjsaif):

Divide both sides by -3. :)

OpenStudy (acacia):

kk

KjSaif (kjsaif):

−3x/−3 = 24/−3

OpenStudy (acacia):

8x?

OpenStudy (acacia):

-8x

KjSaif (kjsaif):

x=−8

KjSaif (kjsaif):

correct! nice job @Acacia! :)

OpenStudy (acacia):

THANKS!

KjSaif (kjsaif):

:) and...................

KjSaif (kjsaif):

\(\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\\\color{white}{.}\\\Huge\sf\color{blue}{~~~~Welcome~to~OpenStudy!~\ddot\smile}\\\color{white}{.}\\\\\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\)